# LLaVA-PruMerge: Adaptive Token Reduction for Efficient Large Multimodal Models

### v4

@@ Abstract
Large Multimodal Models (LMMs) have shown significant reasoning capabilities by connecting a visual encoder and a large language model. LMMs typically use a fixed amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content. Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which increase the number of visual tokens significantly. However, due to the design of the Transformer architecture, computational costs associated with these models tend to increase quadratically with the number of input tokens. To tackle this problem, we explore a token reduction mechanism and find, similar to prior work, that many visual tokens are spatially redundant. Based on this, we propose PruMerge, a novel adaptive visual token reduction approach, which largely reduces the number of visual tokens while maintaining comparable model performance. We first select the unpruned visual tokens based on their similarity to class tokens and spatial tokens. We then cluster the pruned tokens based on key similarity and merge the clustered tokens with the unpruned tokens to supplement their information. Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 18 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities. LLMs are usually high-capacity Transformers pretrained with a large-scale text corpus. Large Multimodal Models (LMMs), inherit the LLMs for text generation, but also leverage a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
LMMs need substantial computation to conduct inference. The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM. For example, the commonly used CLIP visual encoder, ViT-L, only has 0.3B parameters, while the corresponding LLM such as LLaMA [Touvron et al., 2023] or Vicuna [Vicuna, 2023] can have 7B or 13B parameters. As a result, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023]. However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap in visual question-answering and reasoning such as VQAv2 and MM-Bench [Chu et al., 2024]. A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation. The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens. For example, in LLaVA-1.5, 576 visual tokens are appended, leading to high training and inference costs. Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
(a) Main idea of LLaVA-PruMerge.
@@
Figure 1: (a) We prune and merge the visual tokens coming from the vision encoder, while keeping all other procedures of the LMM the same. By reducing the number of visual tokens, our proposed method, PruMerge, significantly reduces the computation cost for text generation in LMMs. (b) The visualizations of the attentive tokens. We design a token reduction method to adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently. More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images. Besides, such attentive tokens are usually located at the regions with dense information.
In our study, we find that such visual tokens are redundant, similar to findings in previous related work Bolya et al. [2023], Liu et al. [2022], and most of the visual tokens can be pruned without largely sacrificing the performance. In particular, we find that the activations are very sparse upon the similarity matrix between the class token and spatial patches, which indicates that only a small amount of the visual tokens are related to the key visual information in the image. Motivated by this, we use this similarity to select important visual tokens. Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune the visual tokens. Moreover, we merge the visual tokens using $k$-nearest neighbor and update the sampled visual tokens via weighted averaging, which further enhances performance. Finally, we optionally finetune the LLM to let the model better adapt to our token deduction design.
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Our work demonstrates the effectiveness of building efficient LMMs from the perspective of visual token reduction.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text. Large Multimodal Models (LMMs) [Liu et al., 2023b, Zhu et al., 2023, Yin et al., 2023, Zhang et al., 2024] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format. More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023]. Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022]. The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a]. In this paper, we reduce the number of visual tokens by leveraging the similarity between the class token and the spatial patch tokens.
@@ 2.2 Token Reduction
The notorious squared complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the sequence length. Sparse attention such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context. Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching. In recent LMMs [Liu et al., 2023b, Zhu et al., 2023], prefix visual tokens serve as a fixed budget for context, which becomes one of the leading factors for their efficiency. In our study, we find that by simply pruning and merging visual tokens based on their similarity, we can achieve comparable performance using less than one tenth of the original tokens.
@@ 3 Method: Token Pru-Merging
In this section, we first review the basic implementation of large mutilmodal models (LMMs), with a particular focus on the visual encoder component (i.e., Vision Transformer). We highlight the direct correlation between the number of visual tokens and the efficiency of LMMs (Sec. 3.1). Next, we present a plug-and-play token reduction method specifically designed for LMMs, called token PruMerge. Our method features two key components: (1) Adaptive Important Token Selection (AITS) via Outlier Detection which adaptively determines the optimal number of visual tokens to retain based on the unique characteristics of the image (Sec. 3.2); and (2) Token Supplement (TS) via Similar Key Clustering, which facilitates more efficient processing without compromising the model’s performance by maintaining the integrity and richness of the visual information (Sec. 3.3).
@@ 3.1 Preliminaries
Figure 2: The conceptual idea of LLaVA-PruMerge. Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens; (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster. Here $m$ denotes the visual token compression ratio.
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for large multimodal models, in which the input image is converted into a sequence of representative tokens by the ViT, and then feed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. An input image is divided into patches and each patch is projected into a token embedding by ViTs. In addition to the set of image tokens, a new class token (i.e., [CLS] token) is added to aggregate global image information and perform classification. Let $n$ denote the number of input tokens (image patches) to the ViT. Let $n$ denote the number of input tokens to the ViT. ViTs consist of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), shortcut connections, and layer normalization [Ba et al., 2016]. These components work together to improve the model’s capability to understand visual data [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. The self-attention computes the relevance of each item to other items:
@@
where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$. In the last layer of ViT, the [CLS] token is used for classification. Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
@@
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, as well as $\sigma$ denoting the nonlinear activation function. The general forward pass of ViT is illustrated in the left part of Figure 2.
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which can map the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of LMMs is depicted in Figure 1. It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022]. There is an increasing demand for processing of high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Figure 3: (a) The distribution of attention values (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token. (b) Image Visualizations of PruMerge and PruMerge+.
The most straight-forward solution to improve the efficiency of visual token utilization in LMMs is to prune the redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022]. To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
@@
Interquartile Range (IQR) Method for outlier detection. To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [Boukerche et al., 2020], a statistical technique known for its robustness in outlier detection. The essence of the IQR method lies in its capability to establish a boundary or “fence” that delineates the normal range of data. This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1. Specifically, the computation is as follows: the “lower fence” is set at 1.5 $\times$ IQR below Q1, and the “upper fence” is set at 1.5 $\times$ IQR above Q3. Any attention values residing outside these fences are classified as outliers. Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context. Note that we use the class attention value from the penultimate layer for this calculation.
@@
As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image. Simpler images such as ”Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens. (2) The sampled tokens are spatially aligned with the important details, while less tokens are assigned for the background. Such visualizations align with our visual token sampling design. These trends are also observed at the benchmark level. In Table 4, the average token numbers on various benchmarks differ.

### v5

@@ Abstract
Large Multimodal Models (LMMs) have shown significant visual reasoning capabilities by connecting a visual encoder and a large language model. LMMs typically take in a fixed and large amount of visual tokens, such as the penultimate layer features in the CLIP visual encoder, as the prefix content. Recent LMMs incorporate more complex visual inputs, such as high-resolution images and videos, which further increases the number of visual tokens significantly. However, due to the inherent design of the Transformer architecture, the computational costs of these models tend to increase quadratically with the number of input tokens. To tackle this problem, we explore a token reduction mechanism that identifies significant spatial redundancy among visual tokens. In response, we propose PruMerge, a novel adaptive visual token reduction strategy that significantly reduces the number of visual tokens without compromising the performance of LMMs. Specifically, to metric the importance of each token, we exploit the sparsity observed in the visual encoder, characterized by the sparse distribution of attention scores between the class token and visual tokens. This sparsity enables us to dynamically select the most crucial visual tokens to retain. Subsequently, we cluster the selected (unpruned) tokens based on their key similarity and merge them with the unpruned tokens, effectively supplementing and enhancing their informational content. Empirically, when applied to LLaVA-1.5 [Liu et al., 2023a], our approach can compress the visual tokens by 14 times on average, and achieve comparable performance across diverse visual question-answering and reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) [OpenAI, 2023b, Team et al., 2023, Jiang et al., 2023, Touvron et al., 2023] have shown strong reasoning abilities. LLMs are usually high-capacity Transformers [Vaswani et al., 2017] pretrained with a large-scale text corpus. Large Multimodal Models (LMMs), inherit LLMs for text generation, while also leveraging a visual encoder such as CLIP-ViT [Radford et al., 2021] to embed image patches into visual tokens as the prefix visual context.
LMMs need substantial computation for inference. The LLM is the primary factor for the high computation cost, since the visual encoder is usually quite small relative to the LLM. For example, the commonly used CLIP visual encoder, ViT-L, only has 0.3B parameters, while the corresponding LLM such as LLaMA [Touvron et al., 2023] or Vicuna [Vicuna, 2023] can have 7B or 13B parameters. As a result, reducing the LLM’s inference cost is the key to achieving low LMM inference cost.
Prior works [Chu et al., 2023, 2024, Yuan et al., 2023a] mainly focus on replacing the LLM backbone with a smaller language model with less parameters, such as Phi-2 [Javaheripi et al., 2023]. However, such approaches sacrifice the reasoning abilities of LLMs, leading to a large performance gap on visual question-answering and reasoning tasks such as VQAv2 and MM-Bench [Chu et al., 2024]. A similar approach is to apply quantization for LLMs [Liu et al., 2023b, Yuan et al., 2024].
However, the cost of LLMs comes from not only its large number of parameters, but also the length of the input context due to the quadratic complexity of the Transformer’s attention operation. The context length in LMMs is especially important, where a fixed amount of visual tokens serves as the prefixed tokens. For example, in LLaVA-1.5, 576 visual tokens are appended, and in Video-LLaVA [Lin et al., 2023] that number is even higher, leading to high training and inference costs. Thus, an intriguing question is: Can we reduce the number of prefix visual tokens while maintaining comparable performance?
(a) Main idea of PurMerge.
@@
Figure 1: (a) We prune and merge visual tokens produced by the vision encoder, while keeping all other procedures of the LMM the same. By reducing the number of visual tokens, PruMerge, significantly reduces the computation cost for text generation in LMMs (around 4-10 times in FLOPs for LMM prefill), while can maintain comparable performance. (b) A visualization of the selected tokens. PruMerge can adaptively select visual tokens based on the information density of the visual input, enabling the LLM to perceive visual input effectively and efficiently. More attentive tokens are sampled in complex images such as ones with text, while fewer are sampled on simpler images. The attentive tokens are usually located at the regions with dense information.
In our study, we find that many visual tokens are redundant, similar to findings in previous related work [Bolya et al., 2023, Liu et al., 2022], and most of the visual tokens can be pruned with little sacrifice in performance. In particular, the similarity (i.e., attention scores in the visual encoder’s self-attention module) between the class token and spatial patches are sparse, indicating that only a small number of visual tokens are related to key visual information in the visual samples. Motivated by this, we use this sparse similarity to adaptively select important visual tokens, as shown in Fig.1(b). Specifically, we leverage the Interquartile Range (IQR) [Boukerche et al., 2020] scoring function in outlier detection to prune unimportant visual tokens. Moreover, we merge the visual tokens using $k$-nearest neighbor and update the selected important visual tokens via weighted averaging, which further enhances performance. Finally, we design PruMerge+, which samples visual tokens spatial-uniformly to complement the unpruned tokens. PruMerge+ not only minimizes performance degradation but also ensures substantial token reduction, maintaining a more comprehensive and representative selection of visual tokens.
Empirically, PruMerge can effectively and adaptively reduce the visual tokens in each image in LLaVA-1.5 [Liu et al., 2023a], where with just 5.5% of visual tokens, which is around 32 tokens for an image on average, LLaVA-PruMerge can maintain comparable performance with that of retaining all 576 tokens across diverse benchmarks. Furthermore, PruMerge showcases its versatility across various modalities, including video. By integrating PruMerge with Video-LLaVA during the inference phase alone—eliminating the need for additional training—we not only expedite processing within video-LLMs but also enhance their performance across multiple benchmarks.
@@ 2.1 Efficient Large Multimodal Models (LMMs)
Large Language Models (LLMs) such as GPT-4 [OpenAI, 2023b], LLaMA [Touvron et al., 2023], Mistral [Jiang et al., 2023], and Gemini [Team et al., 2023] have demonstrated strong question answering and reasoning capabilities over text. Large Multimodal Models (LMMs) [Liu et al., 2023b, Zhu et al., 2023, Yin et al., 2023, Zhang et al., 2024] extend these reasoning capabilities to images, where given an image and an associated question, a vision encoder and an LLM are leveraged to generate text responses in a chat format. More recent works extend whole-image understanding into region-level understanding [Cai et al., 2024, Zhang et al., 2023b, Peng et al., 2023, Chen et al., 2023], video understanding [Lin et al., 2023, Zhang et al., 2023a] and 3D scene understanding [Hong et al., 2023]. Such works typically feed the visual tokens directly into the LLM as prefix tokens, via either an MLP [Liu et al., 2023a], Qformer [Dai et al., 2023, Zhu et al., 2023], or resampler [Alayrac et al., 2022]. The number of visual tokens can be prohibitively long, especially when the images are high-resolution [Liu et al., 2024, OpenAI, 2023a]. In this paper, we reduce the number of visual tokens with a novel adaptive prune and merge procedure.
@@ 2.2 Token Reduction
The notorious quadratic complexity in Transformers [Vaswani et al., 2017] is a well-known problem, as it is one of the key bottlenecks in scaling the input sequence length. Sparse attention methods such as Linformer [Wang et al., 2020] and ReFormer [Kitaev et al., 2020] reduce the quadratic attention complexity by conducting attention operations within a certain region rather than the full context. Token merging [Bolya et al., 2023] utilizes full attention but gradually reduces the number of tokens in each transformer block by selecting the most representative tokens with bipartite matching. One of the main causes of the inefficiency of recent LMMs is their use of a large number of prefix visual tokens that serve as a fixed budget for context [Liu et al., 2023b, Zhu et al., 2023]. In our study, we present a novel plug-and-play token reduction method based on visual token similarities, which achieves comparable performance with less than one tenth of the original tokens.
@@ 3 Method: Token Pru-Merging
In this section, we first review the basic implementation of large mutilmodal models (LMMs), with a particular focus on the visual encoder component (i.e., Vision Transformer). We highlight the direct correlation between the number of visual tokens and the efficiency of LMMs (Sec. 3.1). Next, we present a plug-and-play token reduction method specifically designed for LMMs, called token PruMerge. Our method features two key components: (1) Adaptive Important Token Selection (AITS) via Outlier Detection which adaptively determines the optimal number of visual tokens to retain based on the unique characteristics of the image (Sec. 3.2); and (2) Token Supplement (TS) via Similar Key Clustering, which facilitates efficient processing without compromising the model’s performance by maintaining the integrity and richness of the visual information (Sec. 3.3).
@@ 3.1 Preliminaries
Figure 2: The conceptual idea of PruMerge. Our approach has 3 steps: (1) Sample important tokens according to the similarities between the class tokens and spatial visual tokens via an outlier detection algorithm (see Sec.3.2); (2) Cluster the visual tokens via $k$-nearest neighbor; and (3) Adjust the sampled visual tokens via weighted averaging for each cluster (see Sec.3.3). Here $m$ denotes the visual token compression ratio.
Vision Transformers (ViTs) [Dosovitskiy et al., 2020] are the most widely used vision encoder for LMMs, in which the input image is converted into a sequence of representative tokens by the ViT, and then fed into an LLM for understanding [Liu et al., 2024, Zhu et al., 2023, Hong et al., 2023, Zhang et al., 2024]. An input image is divided into a grid of patches and each patch is projected into a token embedding by the ViT. In addition to the patch tokens, a class token (i.e., [CLS] token) is computed to aggregate global image information for classification. A ViT consists of a set of transformer blocks, which in turn consist of several essential components: a multi-head self-attention (MSA) layer, a feed-forward neural network (FFN), skip connections, and layer normalization [Ba et al., 2016]. These components work together to improve the model’s capability to understand visual data [Han et al., 2022]. In the self-attention layer, an input token is projected into three distinct vectors: query $\mathbf{q}$, key $\mathbf{k}$, and value $\mathbf{v}$, using three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. These vectors, corresponding to different inputs, are assembled into matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, respectively. The self-attention computes the relevance of each item to other items:
@@
where attention matrix $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\mathbf{T}}}{\sqrt{d_{k}}}\right)$ and $d_{k}$ is the dimension of $\mathbf{q}$ and $\mathbf{k}$. In the last layer of the ViT, the [CLS] token is used for classification. Similarly, the attention between [CLS] token and other visual tokens is computed by the attention mechanism:
@@
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the matrices of the linear transformation layers, and $\sigma$ denotes the nonlinear activation function. The general forward pass of ViT is illustrated in the left part of Figure 2.
Large Multimodal Models (LMMs). Following the forward pass through a Vision Transformer (ViT), a set of visual tokens is generated. These tokens are then processed by the input projector $\mathbf{\Theta}_{\mathbf{X}\rightarrow\mathbf{T}}$, which maps the encoded visual features from $\mathbf{F}_{X}$ into the text feature space $\mathbf{T}$. The aligned features and the text prompts $\mathbf{P}_{T}$ are then fed into the LLM backbone [Zhang et al., 2024]. The overall architecture of an LMM is depicted in Figure 1.
It is important to note that the computational cost with these models increases quadratically with the number of input tokens to the LLM [Tay et al., 2022]. Mathematically, if there are $N$ tokens in the input, the self-attention mechanism computes a $N\times N$ matrix of attention scores, where each entry in this matrix represents the attention score between a pair of tokens. However, there is an increasing demand for processing high-resolution images and videos, which increases the number of visual tokens, further exacerbating computation costs. The reduction of visual tokens presents a promising approach to improving the efficiency of LMMs by reducing the escalating computational requirements.
@@ 3.2 Adaptive Important Token Selection via Outlier Detection
Figure 3: (a) The distribution of attention scores (in the penultimate layer of CLIP-ViT) between the [cls] token and visual tokens. The values in the y-axis are logarithmic values. Notably, most of the spatial visual tokens have near-zero attention values with the class token. (b) Visualizations of PruMerge and PruMerge+.
The most straightforward solution to improve the efficiency of visual token utilization in LMMs is to prune redundant visual tokens [Liu et al., 2022, Yin et al., 2022, Liang et al., 2022]. To realize token pruning, we need to address a pivotal question: How do we determine the importance of each visual token?
@@
Interquartile Range (IQR) Method for outlier detection. To identify outliers within class attention values, we adopt the Interquartile Range (IQR) method [Boukerche et al., 2020], a statistical technique known for its robustness in outlier detection. The essence of the IQR method lies in its capability to establish a boundary or “fence” that delineates the normal range of data. This is achieved by calculating the IQR (the difference between the third quartile Q3 and the first quartile Q1) and subsequently defining the outer limits of the normal range as 1.5 times the IQR above Q3 and below Q1. Specifically, the computation is as follows: the “lower fence” is set at 1.5 $\times$ IQR below Q1, and the “upper fence” is set at 1.5 $\times$ IQR above Q3. Any attention values residing outside these fences are classified as outliers. In practice, only the “upper fence” is activated since the attention score is positive. Through this method, we can adaptively identify and select the visual tokens for each image that exhibit outlier attention values, i.e., those playing a significant role in representing the image within the LMM context. Note that we use the class attention value from the penultimate layer for this calculation.
@@
As shown in Figure 1(b), the sampled visual tokens demonstrate two behaviors: (1) The number of attentive tokens are proportional to the complexity of the image. Simpler images such as “Billboard among blue sky” owns fewer tokens while images with rich information such as a screen with dense texts own more tokens. (2) The sampled tokens are typically spatially aligned with important content. Such visualizations align with our visual token sampling design. These trends are also observed at the benchmark level; in Table 6, the average token numbers on various benchmarks differ.
3.3 Token Supplement via Similar Key Clustering
Following the selection of informative visual tokens, we next optimize the utilization of the remaining tokens. While pruned tokens may initially seem extraneous, they hold potential value for the perception capabilities of the LLM backbone. This potential arises particularly in cases where an image contains large object parts that dominate the scene. In such scenarios, overly aggressive pruning could inadvertently diminish the model’s ability to represent the image comprehensively.
To address this, we devise a token merging method aimed at enhancing the representational capacity of the selected unpruned tokens. This method involves the strategic fusion of currently pruned tokens, as depicted in Figure 2. To choose the pruned tokens to merge, we need a way to measure similarity between visual tokens. Here we leverage the self-attention mechanism in ViTs. Since the key vector of each patch token already contains information summarized in the self-attention module [Vaswani et al., 2017], the final layer’s key vector serves as the representation. And then we use the dot product between keys to calculate which tokens have similar visual information [Bolya et al., 2023]:
which yields $\mathbf{K}\mathbf{K}^{T}(i,j)$ for tokens $i,j$ in vectorized form for the set of all tokens ${1,2,\cdots,n}$, where $n$ is the number of input visual tokens.
3.4 PruMerge+: Bridging the Efficiency-Performance Gap
While PruMerge achieves a remarkable reduction in the number of visual tokens—over tenfold compared to the original setup—the process is not without drawbacks. Specifically, the compression technique, though efficient, introduces a marginal performance discrepancy between the original LLaVA model and its PruMerge-optimized counterpart, LLaVA-PruMerge. To address this, we introduce PruMerge+, a refined version that strikes an optimal balance between token reduction and model performance.
PruMerge+ enhances our original method by maintaining the ability to significantly reduce visual token count—by an average of fourfold—with minimal performance degradation. This improvement is detailed in Algorithm 1, building upon the token selection strategies outlined in Section 3.2. A new aspect of PruMerge+ lies in its enhanced token selection process. Beyond merely focusing on the previously identified important tokens, PruMerge+ extends its reach to encompass additional visual tokens from areas initially deemed less critical. This is achieved through a spatially uniform sampling of visual tokens, guided by a predetermined ratio informed by the distribution of outlier tokens. This methodology ensures a more comprehensive and representative selection of visual tokens, as depicted in Figure 3(b), thereby minimizing performance losses while still achieving substantial token reduction.